An auto rickshaw charges Rs 10 for the first kilometre then Rs 8 for each such subsequent kilometre. The total charge (in Rs) for d kilometres is __________.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

An auto rickshaw charges Rs 10 for the first kilometre then Rs 8 for each such subsequent kilometre. The total charge (in Rs) for d kilometres is 2 + 8d.

From the question it is given that,
An auto rickshaw charges₹10 for the first kilometre
Then₹8 for each such subsequent kilometre.
So, the total charge (in Rs) for d kilometres is = 10 + (d - 1)8

= 10 + 8d–8
= 2 + 8d

Thanks